# MambaIO: Global-Coordinate Inertial Odometry for Pedestrians via Multi-Scale Frequency-Decoupled Modeling

### v1

@@ V Conclusion
This work systematically investigates the influence of coordinate frame selection on pedestrian inertial odometry, highlighting the representational advantages of the global coordinate frame in non-rigid carrier scenarios. Building on this insight, we propose MambaIO, which employs a Laplacian pyramid to separate high- and low-frequency components of IMU signals and integrates a multi-path convolution module with a Mamba module to jointly model local details and global motion trends. Despite Mamba’s linear complexity, which is lower than that of conventional Transformers, its inference speed remains relatively limited when processing long sequences, particularly in resource-constrained mobile deployment settings. Future research will focus on developing even lighter and more efficient sequence modeling strategies to further improve real-time performance while maintaining accuracy.

### v2

@@ V Conclusion
This work systematically investigates the impact of coordinate frame selection on pedestrian IO and reveals the advantages of the global frame in motion representation under non-rigid carrier scenarios. Building upon this insight, we propose MambaIO, which decomposes IMU signals into high- and low-frequency components using a Laplacian pyramid and jointly models local details and global motion trends through a multi-path convolution module and a Mamba module. However, the high-frequency component still contains interference noise that is difficult to isolate precisely, and effective separation mechanisms remain lacking—a key direction for future research.